OpenStudy (anonymous):

A triangle has side lengths 10, 15, and 7. Is the triangle acute, obtuse, or right? Explain

OpenStudy (anonymous):

MY ANSWER: Obtuse, because to find if the triangle is right, acute or obtuse you use the Pythagorean Theorem. You plug in the two smaller numbers in the equation to it looks like this, 7^2+10^2=12.2. Since the hypotenuse given in the problem is larger than the hypotenuse you get in the theorem it is obtuse. Can someone explain what I did wrong?

OpenStudy (anonymous):

You're correct.

OpenStudy (anonymous):

But my teacher said it was wrong...

OpenStudy (anonymous):

The answer would be awkward, I know.

OpenStudy (anonymous):

Awkward?

OpenStudy (anonymous):

Lol... tell your teacher to look again.

OpenStudy (anonymous):

I did, and she still said it was wrong. That is why I am asking for help.

OpenStudy (anonymous):

Stop cheating.

OpenStudy (anonymous):

Cheating is so pathetic, seriously.

OpenStudy (anonymous):

Don't cheat, you learn nothing by cheating.

OpenStudy (anonymous):

I'm not?

OpenStudy (anonymous):

@e.mccormick Can you help?

OpenStudy (anonymous):

@e.mccormick Don't help him, he's cheating.

OpenStudy (anonymous):

I'm not cheating, so if you can't help please leave this thread alone.

OpenStudy (e.mccormick):

The only thing I see wrong is ytour math: \(7^2+10^2=12.2\) is not true. \(7^2+10^2=149\) is true and \(15^2=225\) is true and \(\sqrt{7^2+10^2}\approx12.2\) is true.

OpenStudy (anonymous):

Okay, thats it?

OpenStudy (e.mccormick):

If that is the exact answer you submitted, yes, that is why it would be marked wrong. you did not actually apply the Pythagorean Theorem in that equation. You came close, but forgot the root sign!

OpenStudy (anonymous):

Oh okay! Thanks!